# Stoichiometry Representation Learning with Polymorphic Crystal Structures

### v1

@@ 5.2 Empirical Results
Physical Validity of Predicted Properties. To further verify the physical validity of predicted properties, we theoretically calculate the figure of merit $Z\bar{T}$ 333 In thermoelectric materials, the figure of merit $Z\bar{T}$ plays a fundamental role in determining how effectively power can be generated and energy can be harvested across various applications Nozariasbmarz et al. (2020). of thermoelectrical materials with the predicted properties in ESTM datasets in Table 1. More specifically, given predicted electrical conductivity (E.C.) $\sigma$, thermal conductivity (T.C.) $\lambda$, Seebeck coefficient $S$, we can compute the figure of merit $Z\bar{T}$ as follows: $Z\bar{T}=\frac{S^{2}\sigma}{\lambda}\bar{T}$, where $\bar{T}$ indicates a conditioned temperature, i.e., 300 K and 600 K. In Table 1, we have following observations: 1) Looking at the general model performance on ESTM datasets and $Z\bar{T}$, we find that performing well on ESTM datasets does not necessarily indicate the predictions are physically valid. 2) In contrast, models that incorporate structural information tend to produce physically valid predictions in both ESTM datasets, underscoring the importance of the crystal structural information. 3) Moreover, PolySRL consistently outperforms baseline methods, demonstrating that PolySRL not only learns accurate representations of stoichiometry but also ensures the physical validity of the predictions. We provide further analysis on the predicted $Z\bar{T}$, and high throughput screening results of thermoelectrical materials in Appendix F.3.
@@ 5.3 Uncertainty Analysis
Number of Structures. In this section, we examine how uncertainties vary according to the number of possible structures. To do so, we first collect all possible structures of stoichiometry in Band Gap dataset from MP database2 and Open Quantum Materials Database (OQMD) [^4] https://oqmd.org/. Subsequently, we compute the average uncertainties for stoichiometry groups with the same number of possible structures. In Figure 3 (a), we have the following observations: 1) In general, the uncertainty of stoichiometry that has polymorphic structures (##\## possible structures $\geq 2$) was higher than that of the stoichiometry with a single structure (##\## possible structures $=1$), demonstrating that PolySRL learns uncertainties regarding polymorphic structures.
@@
2) On the other hand, an increase in the number of possible structures in OQMD leads to an increase in the uncertainty, demonstrating that PolySRL learns uncertainties related to the diverse polymorphic structures. Note that this trend is mainly shown in the OQMD dataset due to the fact that OQMD encompasses not only realistic but also theoretically possible structures, indicating that PolySRL acquires knowledge of theoretical uncertainties in materials science. 3) Furthermore, we notice high uncertainties when there are no potential structures available (i.e., when ##\## possible structures $=0$) in comparison to stoichiometry with a single possible structure, suggesting that uncertainty contains information about the computational feasibility of the structure.
@@
Impurity of Materials. Next, we investigate how impurities in materials influence the uncertainty in stoichiometry. Specifically, we compare the average stoichiometry uncertainty between groups of doped or alloyed materials (i.e., Impure) and their counterparts (i.e., Pure) in thermoelectric materials datasets, i.e., ESTM 300K and ESTM 600K, where doping and alloying are commonly employed to enhance their performance. In Figure 3 (b), we notice a substantial increase in the uncertainty within impure materials compared with their pure counterparts. This observation is in line with common knowledge in materials science that doping or alloying can lead to chaotic transformations in a conventional structure Kawai et al. (1992); Jin & Nobusada (2014), demonstrating that PolySRL also captures the complexity of structure as the uncertainty. In conclusion, uncertainty analysis highlights that PolySRL effectively captures the uncertainty related to the presence of polymorphic structures within a single stoichiometry and the computational challenges associated with crystal structures.
Figure 4: Case studies.
@@
Case Studies. While our previous analysis on uncertainties generally aligns with our expectations, we do observe some instances where PolySRL exhibits high uncertainty in non-polymorphic stoichiometries and minimal uncertainty in polymorphic stoichiometries. First, we observe the stoichiometry of HgCl and CaS exhibit high uncertainty, even though they only have one possible structure (Figure 4 (a)). We attribute this phenomenon to the limited availability of element combinations in the MP dataset, which occurred due to several factors, including the rarity of certain elements and the difficulty in synthesizing substances with specific combinations of elements Castor et al. (2006); Jang et al. (2020). On the other hand, we observe the learned distribution of ScN and $\text{AgSO}_{4}$ collapsed to zero even though each of them has three possible polymorphic structures (Figure 4 (b)). This behavior arises from the structural similarity among the polymorphic structures, where all three polymorphic structures of each stoichiometry fall within the same cubic and monoclinic structural system, respectively. In conclusion, PolySRL acquires detailed insights concerning polymorphic structures beyond mere quantitative counts. Additionally, we include further analysis on the correlation between uncertainty and model performance, along with supplementary case studies that are in line with our anticipated results in Appendix F.5.
6 Conclusion

### v2

@@ 3.1 Quantitative Results
Physical Validity of Predicted Properties. To further verify the physical validity of predicted properties, we theoretically calculate the figure of merit $Z\bar{T}$ 111 In thermoelectric materials, the figure of merit $Z\bar{T}$ plays a fundamental role in determining how effectively power can be generated and energy can be harvested across various applications [41]. of thermoelectrical materials with the model-predicted properties in ESTM datasets in Table 1. Note that the model-predicted properties are assumed to be physically valid, if the theoretically calculated figure of merit $Z\bar{T}$ is accurate (i.e., low MAE for $Z\bar{T}$ in Table 1). More specifically, given predicted electrical conductivity (E.C.) $\sigma$, thermal conductivity (T.C.) $\lambda$, Seebeck coefficient $S$, we can compute the figure of merit $Z\bar{T}$ as follows: $Z\bar{T}=\frac{S^{2}\sigma}{\lambda}\bar{T}$, where $\bar{T}$ indicates a conditioned temperature, i.e., 300 K and 600 K. In Table 1, we have following observations: 1) Looking at the general model performance on ESTM datasets and $Z\bar{T}$, we find that performing well on ESTM datasets does not necessarily indicate the predictions are physically valid. For instance, while training the model on DFT-calculated bandgap (MP Band G.) consistently yields more accurate predictions on the ESTM dataset compared to GraphCL, the $Z\bar{T}$ values derived from MP Band G. predictions are significantly worse than those obtained from GraphCL. 2) Overall, models that incorporate structural information tend to produce physically valid predictions in both ESTM datasets, underscoring the importance of the crystal structural information. 3) Moreover, PCRL consistently outperforms baseline methods, demonstrating that PCRL not only learns accurate representations of composition but also ensures the physical validity of the predictions.
Figure 2: Scatter plot between true and predicted $Z\bar{T}$.
Figure 3: High-throughput screening results.
Further Analysis. In this section, we delve deeper into the physical validity of predicted properties for thermoelectrical materials by observing scatter plots that compare the actual ground truth values of $Z\bar{T}$ with the values obtained by the model predictions. For clearer visualization, we select one baseline model from models that consider DFT-calculated properties (i.e., MP Band G.) and structures (i.e., 3D Infomax). In Figure 2, we notice that the predictions produced by PCRL consistently yield accurate calculations of $Z\bar{T}$ without any outliers. This observation underscores the model’s ability to predict physically valid properties for thermoelectrical materials. Additionally, we observe that the model, specifically MP Band G., which lacks consideration of the structural information within composition, tends to produce outliers more frequently when contrasted with models that incorporate structural information. More specifically, three outliers made by MP Band G. in Figure 2 (a) are $\text{Co}_{9}\text{S}_{8}$, $\text{Cu}_{5}\text{Sn}_{2}\text{S}_{6.65}\text{Cl}_{0.35}$, and $\text{Cu}_{5.133}\text{Sn}_{1.866}\text{S}_{6.65}\text{Cl}_{0.35}$. In case of $\text{Co}_{9}\text{S}_{8}$, there exist only one possible structure in MP dataset, and there was no existing structure for $\text{Cu}_{5}\text{Sn}_{2}\text{S}_{6.65}\text{Cl}_{0.35}$, and $\text{Cu}_{5.133}\text{Sn}_{1.866}\text{S}_{6.65}\text{Cl}_{0.35}$. This suggests that MP Band G. encounters difficulty in acquiring accurate physical properties for materials where obtaining structural information is computationally challenging. On the other hand, in Figure 2 (b), two outliers made by MP Band G. are GeTe and SnTe, each of which has three possible structures in MP dataset. This indicates that MP Band G. suffers from obtaining valid physical properties from polymorphic structures. In conclusion, we argue that this finding underscores the significance of incorporating structural information for accurate predictions.
High-Throughput Screening. As described in the main manuscript, the figure of merit $Z\bar{T}$ determines how effectively power can be generated and energy can be harvested across various real-world applications. To discover novel materials of high $Z\bar{T}$, we perform high-throughput screening based on the predicted $Z\bar{T}$ in Figure 3. In particular, for thermoelectrical materials at room temperature (300 K), we establish a threshold of $Z\bar{T}=0.8$, and for high-temperature scenarios (600 K), we use a threshold of $Z\bar{T}=1.1$. We observe that PCRL outperforms all other baseline methods in ESTM 300K datasets while performing competitively with 3D Infomax in ESTM 600K. This again demonstrates the importance of structural information in compositional representation learning, which has been overlooked in previous works [9, 24].
Model DFT Poly. Castelli Refractive Shear Bulk Exfoliation MP
Prop. Str. Perovskites Index Modulus Modulus Energy Band G. Form. E.
Rand init. 0.140 0.394 0.115 0.085 54.49 0.354 0.119
(0.004) (0.091) (0.003) (0.003) (16.08) (0.005) (0.002)
GraphCL 0.145 0.386 0.117 0.084 56.20 0.351 0.121
(0.006) (0.094) (0.002) (0.002) (17.51) (0.004) (0.001)
MP Band G. 0.141 0.399 0.116 0.085 56.14 0.354 0.119
(0.004) (0.085) (0.002) (0.004) (16.86) (0.007) (0.002)
MP Form. E. 0.134 0.379 0.108 0.080 55.40 0.338 0.115
(0.004) (0.093) (0.002) (0.003) (17.29) (0.002) (0.001)
3D Infomax 0.147 0.388 0.117 0.088 54.47 0.354 0.116
(0.004) (0.094) (0.003) (0.004) (15.24) (0.005) (0.002)
PCRL 0.132 0.394 0.107 0.083 52.90 0.328 0.112
(0.007) (0.092) (0.002) (0.003) (14.00) (0.005) (0.003)
Model Performance in DFT-calculated properties. Although DFT-calculated properties frequently differ from actual wet-lab experimental properties [42], we have included experimental outcomes for seven DFT-calculated properties from the Matbench dataset [43]. These Matbench datasets were assessed using a five-fold cross-validation approach with train/validation/test splits set at a ratio of 72/8/20, as given in previous work [24]. In Table 2, we have following observations: 1) In the DFT-based dataset, we observed significant disparities in trends compared to the experimental datasets in Table 1, demonstrating the inherent difference between the experimental data and DFT-calculated data. For instance, we noticed that the MP Form. E. model consistently outperforms the MP Band G. and 3D Infomax models. 2) Furthermore, given that the datasets are designed to pick the target value linked to the lowest formation enthalpy among different polymorphic structures for a single composition, we find that models trained with specific DFT-calculated values (i.e., Prop. ) do not outperform models trained on corresponding datasets. This discrepancy is attributed to properties derived from non-lowest formation enthalpy polymorphic structures, which can introduce confusion to the model. 3) However, we observe PCRL generally outperforms baseline models, demonstrating its effectiveness in not only wet-lab experimental datasets but also in DFT-calculated datasets.
Fine-Tuning. In this section, we compare the models’ performance in fine-tuning scenarios in Table 3. It is worth noting that the transfer of knowledge from the pre-training phase to downstream tasks, which have limited labeled data, via fine-tuning is vital in materials science due to the common data scarcity problem in wet-lab experimental data. We have the following observations: 1) Although recently proposed supervised learning approaches (i.e., Roost and CrabNet) are specifically tailored for material property prediction solely based on compositional data with a large number of parameters, we observe randomly initialized variant of PCRL (i.e., Rand init.) outperforms the methods. This is because the previous works require an extensive dataset computed via DFT for model training, limiting their applicability in wet-lab experimental data, which is more realistic in real-world materials discovery. 2) From the comparisons between the randomly initialized variant of PCRL (i.e., Rand init.) and the baseline methods in Table 3, we observe that the pre-training step sometimes detrimentally affects model performance, i.e., negative transfer. This indicates that without an elaborate design of the tasks, pre-training may incur negative knowledge transfer to the downstream tasks [44]. 3) However, by comparing Rand init. with PCRL in Table 3, we observe that pre-training the model with PCRL objective consistently leads to performance gain in various downstream tasks, i.e., positive transfer. We attribute this to the probabilistic representation, which maintains a high variance for uncertain materials, thereby preventing the representations of the materials from overfitting to the pretraining task. Given the widespread issue of data scarcity in wet-lab experimental data, this highlights the practicality of PCRL within the materials discovery process. In Appendix C.3, we assess the transferability of various pre-training strategies in an out-of-distribution context, where the datasets are divided according to the types of elements involved.
Table 3: Fine-tuning performance (measured by MAE).
Model Band G. Form. E. Metallic ESTM 300 K ESTM 600 K $Z\bar{T}$
E.C. T.C. Seebeck E.C. T.C. Seebeck 300K 600K
Pre-training
Rand init. 0.390 0.599 0.204 0.849 0.202 0.425 0.659 0.209 0.402 0.073 0.218
(0.012) (0.053) (0.014) (0.174) (0.027) (0.048) (0.098) (0.019) (0.082) (0.024) (0.076)
Roost 0.384 0.743 0.199 0.851 0.216 0.406 0.684 0.240 0.402 0.066 0.297
(0.008) (0.069) (0.023) (0.126) (0.037) (0.046) (0.180) (0.048) (0.054) (0.005) (0.031)
CrabNet 0.403 0.759 0.220 1.016 0.285 0.491 0.816 0.309 0.691 0.140 0.574
(0.008) (0.052) (0.017) (0.153) (0.049) (0.088) (0.167) (0.023) (0.057) (0.063) (0.347)
Pre-training
GraphCL 0.391 0.607 0.193 0.862 0.198 0.412 0.643 0.205 0.412 0.064 0.202
(0.011) (0.026) (0.018) (0.236) (0.031) (0.006) (0.098) (0.021) (0.077) (0.014) (0.086)
MP Band G. 0.382 0.604 0.193 0.829 0.210 0.405 0.632 0.197 0.402 0.067 0.206
(0.012) (0.036) (0.025) (0.187) (0.038) (0.006) (0.095) (0.028) (0.081) (0.011) (0.119)
MP Form. E. 0.391 0.582 0.197 0.822 0.195 0.410 0.641 0.209 0.428 0.069 0.223
(0.013) (0.015) (0.019) (0.167) (0.031) (0.041) (0.102) (0.043) (0.086) (0.010) (0.097)
3D Infomax 0.391 0.606 0.194 0.844 0.210 0.402 0.633 0.207 0.391 0.066 0.154
(0.006) (0.027) (0.019) (0.195) (0.032) (0.005) (0.133) (0.018) (0.077) (0.010) (0.032)
PCRL 0.386 0.576 0.191 0.822 0.189 0.386 0.626 0.195 0.390 0.058 0.179
(0.021) (0.042) (0.024) (0.162) (0.037) (0.069) (0.161) (0.015) (0.077) (0.012) (0.052)
Figure 4: Sensitivity Analysis on $\beta$.
Effect of the Hyperparameter $\beta$. In this section, we verify the empirical effect of the hyperparameter $\beta$, which controls the weight of the KL divergence loss computed between the learned distributions and the standard normal distribution, in Equation 6. We have the following observations from Figure 4: 1) As the hyperparameter $\beta$ increases, the average variance of the learned distributions (i.e., uncertainty) also increases, and the dimension of the variance vectors that collapse to zero (i.e., collapsed ratio) decreases. This indicates that the KL divergence loss effectively prevents the distributions from collapsing. 2) On the other hand, the performance of PCRL deteriorates as $\beta$ increases, indicating that emphasizing the KL divergence loss too much causes PCRL to struggle in learning high-quality compositional representations. However, reducing $\beta$ does not always result in improved performance, as collapsed distribution may not effectively capture information from polymorphic structures. Hence, selecting an appropriate value of $\beta$ is vital for learning high-quality compositional representations while maintaining a suitable level of uncertainty. This selection process could be a potential limitation, as it requires a trial-and-error approach to determine the optimal value. We provide results of additional ablation studies and analysis on hyperparameters in Appendix C.4.
Table 4: Representation learning performance (MAE) comparison trained with different datasets.
Dataset Band G. Form. E. Metallic MP 0.407 0.592 0.194 (Ours) (0.013) (0.039) (0.017) OQMD 0.425 0.663 0.194 (0.011) (0.035) (0.028)
Pre-training with Other Databases. While our model is initially pre-trained using the MP database, we also explore whether other databases can enhance its performance. To do so, we pre-train PCRL with another database, i.e., Open Quantum Materials Database (OQMD) [^2] https://oqmd.org/, which consists of 640,964 unique compositions and their corresponding 1,222,097 structures. It is widely known that the MP database primarily comprises crystal structures that exist in the real world, whereas the OQMD database includes both existing crystal structures and theoretically possible structures. In Table 3.1, we observe that training with the OQMD database deteriorates model performance, suggesting that training with theoretically possible structures may confuse the model when applied to wet-lab experimental datasets that comprise real-world materials. Consequently, we advocate for pre-training the model with the MP database as a beneficial strategy for the real-world material discovery process. We provide full experimental results on various downstream datasets in Appendix C.5.
Table 5: Representation learning performance (MAE) comparison trained with different relationship.
Relationship Band G. Form. E. Metallic
One-to-Many 0.407 0.592 0.194
(Ours) (0.013) (0.039) (0.017)
Many-to-Many 0.410 0.617 0.210
(0.012) (0.025) (0.023)
Many-to-many Relationship between Composition and Structures. In material science, polymorphism refers to the phenomenon where a single composition can be related to multiple crystal structures. On the other hand, a single structure can also be related to multiple compositions; different materials may share the same atomic arrangement, known as isostructural [45], resulting in a many-to-many relationship between composition and structure. To this end, we examine whether modeling many-to-many relationships would be beneficial by learning the probabilistic representation of crystal structure. In Table 3.1, we observe that modeling many-to-many relationships between the composition and structure deteriorates model performance. We attribute this to the wet-lab experimental environment, where uncertainties related to isostructural materials are absent due to comprehensive knowledge of the material’s composition. Consequently, we contend that addressing the ambiguities arising from polymorphism, rather than isostructural issues, is more vital for real-world material discovery. We also provide full experimental results on various downstream datasets in Appendix C.6.
@@ 3.2 Uncertainty Analysis
Table 6: Qualitative uncertainty analysis.
Model Accuracy. First, we analyze the correlation between the accuracy of model predictions and the uncertainties present in the Band Gap dataset. To achieve this, we initially categorize composition based on MAE into intervals such as 0.0 to 1.0, 1.0 to 2.0, $\cdots$, and 4.0 to 5.0. For example, Group 1 in Figure 5 (a) contains the set of compositions whose MAEs are in the range from 0.0 to 1.0. We then calculate the average uncertainties of the model for each group. In Figure 5 (a), uncertainty increases as the MAE values increase, demonstrating that the model effectively estimates uncertainties associated with the prediction.
Number of Potential Structures. Moreover, we examine how uncertainties vary according to the number of potential structures. To do so, we first collect all possible structures of composition in Band Gap dataset from MP database  [^3] https://next-gen.materialsproject.org/, and OQMD 2. Subsequently, we compute the average uncertainties for compositional groups with the same number of possible structures. In Figure 5 (b), we have the following observations: 1) In general, the uncertainty of a composition that has polymorphic structures (##\## possible structures $\geq 2$) is higher than that of the composition with a single structure (##\## possible structures $=1$), demonstrating that PCRL learns uncertainties regarding polymorphic structures. 2) Moreover, an increase in the number of possible structures generally leads to an increase in the uncertainty, demonstrating that PCRL learns uncertainties related to the diverse polymorphic structures. Note that this trend is mainly shown in the OQMD dataset, encompassing not only realistic but also theoretically possible structures, indicating that PCRL acquires knowledge of theoretical uncertainties in materials science. We provide case studies demonstrating that PCRL acquires detailed insights concerning polymorphic structures in Section 3.3.
@@
Impurity of Materials. Next, we investigate how impurities in materials influence the uncertainty in composition. Specifically, we compare the average composition uncertainty between groups of doped or alloyed materials (i.e., Impure) and their counterparts (i.e., Pure) in thermoelectric materials datasets, i.e., ESTM 300K and ESTM 600K, where doping and alloying are commonly employed to enhance the thermoelectric performance of materials. In Figure 5 (c), we notice a substantial increase in the uncertainty within impure materials compared with their pure counterparts. This observation is in line with common knowledge in materials science that doping or alloying can lead to chaotic transformations in a stable structure [6, 7], demonstrating that PCRL also captures the complexity of structure as the uncertainty.
Materials Discovery with Uncertainties. In materials science, the presence of many possible structures for a given composition indicates that researchers have a higher degree of freedom to manipulate the substance, allowing for a wider range of applications for the material. By leveraging the uncertainties identified by PCRL, it becomes possible to ascertain if a given composition might present multiple potential structures, thus opening up a broad spectrum of applications without relying on expensive DFT calculations. In this section, we study the materials for which our model provides high uncertainty, specifically focusing on those with five polymorphic structures. Cesium Iodide (CsI) is a material that has the highest uncertainty among the materials that have five polymorphic structures. Notably, CsI is known for its high scintillation efficiency, meaning that it can effectively convert high-energy radiation (e.g., X-rays or gamma rays) into visible light [46]. This characteristic makes CsI widely applicable as a fluorescent agent in areas like chemical analysis, physical experimentation, and medical fields [47, 48, 49]. Lead Telluride (PbTe) ranks as a material with the second-highest uncertainty among those with five polymorphic structures. PbTe is widely used in thermoelectric devices, such as generators and coolers, due to its high thermoelectric efficiency [50, 51]. Additionally, as a semiconductor with a narrow band gap, PbTe is highly responsive to infrared (IR) light, making it a valuable component in infrared detectors and sensors [52, 53].
In conclusion, uncertainty analysis highlights that PCRL effectively captures the uncertainty related to the model prediction, the presence of polymorphic structures within a single composition, and the computational challenges associated with crystal structures, all of which are a common interest in materials science. Moreover, we observe that the materials exhibiting high uncertainty often find diverse applications across numerous scientific domains, highlighting the practical effectiveness of PCRL in materials discovery.
3.3 Qualitative Analysis
@@
While our previous analysis in Section 3.2 shows an increase in the number of possible structures generally leads to an increase in the uncertainty, we do observe some instances where PCRL exhibits high uncertainty in non-polymorphic compositions and minimal uncertainty in polymorphic compositions. First, in Figure 6 (a), we observe that the composition of HgCl and CaS exhibit high uncertainty, even though they only have one possible structure. We attribute this phenomenon to the limited availability of element combinations of Hg, Cl and Ca, S in the MP dataset, which occurred due to several factors including the rarity of certain elements and the difficulty in synthesizing substances with specific combinations of elements [54, 55]. On the other hand, we observe the learned distribution of ScN and $\text{AgSO}_{4}$ collapsed to zero even though each of them has three possible polymorphic structures (Figure 6 (b)). This behavior arises from the structural similarity among the polymorphic structures, where all three polymorphic structures of each composition fall within the same cubic and monoclinic structural system, respectively. That is, while we observe that an increase in the number of polymorphic structures generally leads to an increase in the uncertainty in Figure 5, this observation highlights that PCRL acquires detailed insights concerning polymorphic structures beyond mere quantitative counts.
4 Discussion
4.1 Application Areas
The practical applications of Probabilistic Compositional Representation Learning(PCRL) span a wide range of industries and research areas within materials science. By addressing the limitations posed by detailed structural data requirements and polymorphism, PCRL offers a practical solution that can be leveraged in various contexts. For example, in catalyst development, PCRL can accelerate the discovery of new catalysts by predicting the activity and selectivity of materials based on their elemental composition alone. This can help in identifying promising candidates for catalytic reactions without the need for extensive experimental validation of each structure. In the development of batteries, PCRL can assist in predicting the electrochemical properties of electrode materials, such as lithium-ion batteries or solid-state batteries. This can streamline the process of finding materials with high energy density and stability. For semiconductor devices, selecting materials with optimal electronic properties is crucial. PCRL can help predict the electrical conductivity, bandgap, and other properties of materials based on their composition, aiding in the selection of materials for next-generation electronics. In the pharmaceutical industry, PCRL can be used to predict the solubility, stability, and bioavailability of drug compounds based on their elemental composition. This can speed up the drug discovery process by identifying viable drug candidates early in the pipeline.
4.2 Broad Impact
The introduction of Probabilistic Composition Representation Learning (PCRL) represents a significant advancement in the application of machine learning (ML) to materials science. The broad impact of this work can be understood through several key dimensions:
First, the ability of PCRL to learn compositional representations of material opens up new avenues for material discovery. Once the compositional encoder is pre-trained, it can be used to predict any properties of interest using only compositional information, without the need for costly structural calculations. By expanding the range of materials and properties that can be studied using ML techniques, PCRL has the potential to democratize access to ML-driven insights in materials science, especially for researchers who may lack the resources or expertise to obtain detailed structural data.
Moreover, as demonstrated in Section 3.1, the learned representations in PCRL are highly transferable to a variety of downstream tasks. This high degree of transferability sets PCRL apart from previous compositional representation learning approaches that typically learn target-specific representations. This characteristic is particularly advantageous in the field of materials science, where wet-lab experimental data are often scarce due to the high costs associated with experiments. Consequently, PCRL can be successfully applied to predict properties even in cases with limited data, extending its utility and impact in materials research.
Finally, PCRL offers a novel approach to material discovery by capturing the uncertainties associated with polymorphism. Polymorphism is a common phenomenon in materials science, where a single composition can exist in multiple structural forms. By capturing the diversity of these forms probabilistically, PCRL provides a deep understanding of material behavior that can guide the design and synthesis of new materials, which has been extensively demonstrated in Section 3.2. This is particularly relevant for applications such as catalyst design, battery materials, and pharmaceuticals, where subtle changes in structure can have dramatic effects on performance.
4.3 Theoretical Uncertainty Analysis
Uncertainties in ML models can be categorized into aleatoric uncertainty and epistemic uncertainty. Aleatoric uncertainty refers to the inherent variability or randomness present in the data generation process, whereas epistemic uncertainty pertains to a lack of knowledge or information about the best model or the true underlying process [56, 57, 58]. In this section, we analyze uncertainty within the framework of aleatoric and epistemic uncertainty.
Aleatoric uncertainty. Crystals with identical chemical compositions can exist in various structural forms, known as polymorphs, under different physical conditions such as temperature, pressure, and synthesis methods. These different structural forms introduce inherent uncertainty into the composition data itself. Additionally, measurement errors or fluctuations that occur during the measurement or synthesizing of crystal structures contribute to aleatoric uncertainty.
Epistemic uncertainty. While we model the probability distribution of composition as a Gaussian distribution, the particles that constitute a material are distributed discretely [59]. Consequently, the material itself follows a distinct probability distribution, which is a composite of the discrete distributions of its constituent particles [60]. In other words, the simplicity of the Gaussian distribution introduces epistemic uncertainty by not accurately replicating the distribution of polymorphs. Additionally, most databases are highly biased due to the costly data collection process [61], which means the database used may not capture the entire distribution of polymorphs. This limitation can also lead the model to exhibit high epistemic uncertainty in regions where there is insufficient data for a specific polymorph.
4.4 Related Works
Compositional Representation Learning. Recently, there has been significant research into graphical representations of materials, which often necessitate costly and sometimes impractical DFT calculations of atomic positions. This constraint hinders their use in real-world materials discovery, where frequent atomic-level reorganization necessitates structural calculations throughout the process. Fortunately, material representations can be alternatively constructed solely based on composition, which indicates the concentration of the constituent elements, without any knowledge of the crystal structure [10]. While composition has historically played a role in effective materials design [62, 63], it has been recently demonstrated that deep neural networks (DNNs) tend to outperform conventional approaches when large datasets are available. Specifically, ElemNet [8] takes elemental compositions as inputs and trains DNNs with extensive high-throughput OQMD dataset [64], showing improvements in performance as the network depth increases, up to a point where it reaches 17 layers. Roost [9] utilizes GNNs for compositional representation learning by creating a fully connected graph in which nodes represent elements, allowing for the modeling of interactions between these elements. Instead of the message-passing scheme, CrabNet [24] introduces a self-attention mechanism to adaptively learn the representation of individual elements based on their chemical environment. While these methods are trained for a specific task, PCRL learn generalized compositional representations that are universally applicable for various tasks considering 1) the structural information and 2) polymorphism in crystal, both of which have not been explored before.
Machine Learning for Materials. Recently, ML approaches have become game changers in the field of materials science, where traditional research has heavily relied on theory, experimentation, and computer simulation, which is costly [65, 66, 1]. Among various ML methods, graph neural networks (GNNs) have been rapidly adopted by modeling crystal structures as graphical descriptions inspired by the recent success of GNNs in biochemistry [67, 68, 69]. Specifically, CGCNN [3] first proposes a message-passing framework based on a multi-edge graph to capture interactions across cell boundaries, resulting in highly accurate prediction for eight distinct material properties. Building upon this multi-edge graph foundation, MEGNet [4] predicts various crystal properties by incorporating a physically intuitive strategy to unify multiple GNN models. Moreover, ALIGNN [70] proposes to utilize a line graph, in addition to a multi-edge graph, to model additional structural features such as bond angles and local geometric distortions. Despite the recent success of graph-based approaches, their major restriction is the requirement of atomic positions, which are typically determined through computationally intensive and sometimes infeasible DFT calculations. As a result, their effectiveness is mainly demonstrated in predicting properties for systems that have already undergone significant computational effort, restricting their utility in the materials discovery workflow [10]. We provide additional related works regarding the probabilistic representation learning and crystal structure prediction in Appendix D.
4.5 Limitations and Future Works
Pre-Training Dataset. While we train the model using the MP database, it is important to note that the structures in the MP database predominantly originate from experimental results accumulated over several decades, encompassing a wide range of varying conditions. MP aggregates these structures and relaxes them through energy minimization at T=0 and P=0. This relaxation can either result in significant structural transformations into a more stable polymorph at T=0 and P=0 or cause minor adjustments if the initial experimental structure is already near a local minimum.
Although the polymorphs used in our work may not precisely represent conditions at varying temperatures and pressures, we believe they still capture the relevant physical trends, as the structural data is consistently computed under standard conditions of 0K and 0 atm. Furthermore, numerous prior studies have shown the effectiveness of pretraining models on DFT-calculated properties at 0K and 0 atm for predicting a wide range of experimental properties observed in wet lab environments [42]. For future work, it would be valuable to explore the collection of initial crystal structures under consistent conditions, though this will require additional effort.
Appropriateness of Gaussian Assumption. In this paper, we choose the Gaussian distribution as a prior, which has multiple advantages when incorporated with deep neural networks as follows:
  • Efficient gradient computation is available with a reparameterization trick.
  • Analytical computation of KL divergence is available and theoretically guaranteed.
In fact, the particles that make up a material are distributed discretely [59], and the material itself manifests following a distinct probability distribution, which is a composite of the discrete distributions attributed to its constituent particles [60]. While parameterizing such discrete distributions for each material is nearly impossible, it would be a promising direction for future work to explore more complex prior distributions that can be parameterized by neural networks.
4.6 Conclusion